CVP Analysis of Multiple Products

Steinberg Company produces commercial printers. One is the regular model, a basic model that is designed to copy and print in black and white. Another model, the deluxe model, is a color printer-scanner-copier. For the coming year, Steinberg expects to sell 100,000 regular models and 20,000 deluxe models. A segmented income statement for the two products is as follows:

Regular Model Deluxe Model Total
Sales $15,000,000   $13,600,000   $28,600,000  
Less: Variable costs 9,000,000   8,160,000   17,160,000  
   Contribution margin $6,000,000   $5,440,000   $11,440,000  
Less: Direct fixed costs 1,200,000   960,000   2,160,000  
   Segment margin $4,800,000   $4,480,000   $9,280,000  
Less: Common fixed costs 1,500,800  
   Operating income $7,779,200  

Required:

1. Compute the number of regular models and deluxe models that must be sold to break even. Round your answers to the nearest whole unit.

Regular models units
Deluxe models units

2. Using information only from the total column of the income statement, compute the sales revenue that must be generated for the company to break even. Round the contribution margin ratio to four decimal places. Use the rounded value in the subsequent computation. (Express as a decimal-based amount rather than a whole percentage.) Round the amount of revenue to the nearest dollar.

Contribution margin ratio
Revenue $
Answer #1
1
Regular Model 32,000 units
Deluxe Model 6,400 units
2
Contribution Margin ratio 0.4
Revenue $ 9,152,000
Explanation   1 :-
Regular Model Deluxe Model Total
Sales $ 15,000,000 $ 13,600,000 $ 28,600,000
Less : Variable Cost 9,000,000 8,160,000 17,160,000
Contribution Margin $ 6,000,000 $ 5,440,000 $ 11,440,000
Sales in units 100,000 20,000
Contribution Margin - per unit $ 60 $ 272
Sales Mix 5 1
Weighted Contribution $ 300 $ 272 $ 572
Fixed Costs $ 3,660,800
Break even Point 6,400
Break even point - in units                           32,000                              6,400
* Contribution Margin - per unit = Contribution Margin / Number of units
Regular Model = $ 6,000,000 / 100,000 units = $ 60
Deluxe Model = $ 5,440,000 / 20,000 units = $ 272
**Sales Mix = Regular Model units : Deluxe Model units
= 100,000 : 20,000
= 5 : 1
Weighted Contribution = Contribution Margin per unit * Sales Mix
Total Fixed Costs = Direct Fixed Cost of Regular Model + Direct Fixed Cost of Deluxe Model + Common Fixed Costs
= $ 1,200,000 + $ 960,000 + $ 1,500,800
= 3,660,800
Break even Point = Fixed Costs / Weighted Contribution
= $ 3,660,800 / $ 572
=6,400 units
Break even point - in units
Regular Model = 6,400 * 5 = 32,000 units
Deluxe Model = 6,400 *1 = 6,400 units
Explanation   2 :-
Regular Model Deluxe Model Total
Sales $ 15,000,000 $ 13,600,000 $ 28,600,000
Less : Variable Cost 9,000,000 8,160,000 17,160,000
Contribution Margin $ 6,000,000 $ 5,440,000 $ 11,440,000
Contribution Margin Ratio 0.4
Fixed Costs $ 3,660,800
Break even point in dollars $ 9,152,000
Contribution Margin Ratio = Contribution Margin / Sales
= $ 11,440,000 / $ 28,600,000
= 0.40
Total Fixed Costs = Direct Fixed Cost of Regular Model + Direct Fixed Cost of Deluxe Model + Common Fixed Costs
= $ 1,200,000 + $ 960,000 + $ 1,500,800
= 3,660,800
Break even point in dollars = Fixed Costs / Contribution Margin Ratio
= $ 3,660,800 / 0.4
= $ 9,152,000